AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 294     Entry time: Sun Feb 27 03:49:36 2022
Author: David, Pierre-Michell 
Subject: Current status 

04:06 85K -->120K : 04:13 120K -->95K

current ring 8e7 p/s

Merger, Tape server (WS3): 1600 kB/s, writing/refreshing ok

Leakage current (WS4): 3.26 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-02-27_04-53-12xytot.png  176 kB  | Hide | Hide all
2022-02-27_04-53-12xytot.png
Attachment 2: 2022-02-27_04-52-51exrst.png  156 kB  | Hide | Hide all
2022-02-27_04-52-51exrst.png
Attachment 3: 2022-02-27_04-52-31extot.png  165 kB  | Hide | Hide all
2022-02-27_04-52-31extot.png
ELOG V3.1.4-unknown